# TEOChat: A Large Vision-Language Assistant for Temporal Earth Observation Data

### v1

@@ 6 Conclusion
A Training Datasets
TEOChatlas consists of GeoChat (Kuckreja et al., 2023), an instruction-following dataset with single earth observation images, as well as four temporal earth observation benchmarks (fMoW (Christie et al., 2018), xBD (Gupta et al., 2019), S2Looking (Shen et al., 2021), and QFabric (Verma et al., 2021)). We describe how we process each dataset in more detail below and a summary of the dataset statistics is provided in Table 7. We note that for all datasets, we (1) resize the shorter side to 224 then crop to 224x224 before inputting the images into the network and (2) randomly sample between all instruction-following tasks when creating examples, sometimes randomly sampling between different phrasings of the same task to introduce diversity in the wording. A spatial map of the locations of the examples in TEOChatlas is shown in Figure 5 and a comparison of TEOChatlas to other instruction-following EO datasets is shown in Table 6.
GeoChat  The GeoChat dataset (Kuckreja et al., 2023) is an instruction-following dataset composed of multiple single image EO datasets spanning tasks including object detection derived from SAMRS (which is itself a composition of DOTA (Xia et al., 2018), DIOR (Li et al., 2020), and FAIR1M (Sun et al., 2022)), visual question answering from LRBEN (Lobry et al., 2020) and Floodnet (Rahnemoonfar et al., 2021), and scene classification from NWPU-RESISC-45 (Cheng et al., 2017). The instruction-following tasks in the GeoChat dataset include scene classification (classify the image into one of several classes provided in the prompt), visual question answering (answer a question about the image), referring expression (identify the location of an object given an expression describing its characteristics), region captioning (describe a location represented as a bounding box in the prompt), detailed description (describing the whole image in detail), grounded description (describing the whole image in detail with bounding boxes included in the description), and multi-round conversation (multiple turns of conversation). In total, the GeoChat training dataset consists of 308,861 single instruction-following examples on 106,747 images.
@@ A Training Datasets
B Temporal Instruction-Following Tasks
@@
C Single Image Datasets
@@
D System Prompt
We use the following prompt for the tasks in TEOChatlas.
@@ D System Prompt
[ backgroundcolor=gray!10, linecolor=white, skipabove=10pt, skipbelow=10pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=10pt, innerbottommargin=10pt ]lightgreybox
@@
A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. USER: This is a sequence of {high resolution, optical} satellite images {from ¡sensor¿}: ¡video¿.
where we randomly inject the information in curly brackets when metadata (resolution or sensor) is available and replace the “¡video¿” token with the “¡image¿” token $N$ times, with $N$ the number of images in the sequence. After this prompt, we add the task instruction then “ASSISTANT:”.
E Additional Training Details
@@
F GeoChat Temporal Post-processing
@@
F.1 Temporal Scene Classification
@@ F.2 Change Detection
F.3 Spatial Referring Expression
@@
S2Looking  The spatial referring expression tasks for S2Looking are to identify the constructed or destructed buildings. For both, we prompt GeoChat to identify all buildings in the first image and in the second image, construct a segmentation mask for each, then take a difference. We use the pixels in the first image but not the second as destructed and the pixels in the second but not in the first as constructed.
F.4 Change Question Answering
@@
F.5 Region-based Change Question Answering
@@
QFabric  These tasks are standard classification tasks, so we follow a similar procedure as for temporal scene classification and feed each image in the sequence (either 2 or 5 images) along with the bounding box in the prompt into GeoChat, then take a majority vote across the images to get the final classification.
G Evaluation Setup
Change Detection and Spatial Referring Expression  To evaluate the performance of TEOChat on building change detection (CD) and spatial referring expression (SRE), we follow the standard performance metrics of the building localization (Loc.) and damage classification (Dmg Cls.) subtasks on xBD and building change detection (Det.) subtask on S2Looking. Specifically, for building localization, we convert the bounding box predictions and ground truth polygons to segmentation masks and compute IoU between the predicted and ground truth segmentation masks. For building change detection, we do the same thing but compute per-pixel F1 between the masks. For building damage classification, we input the ground truth polygon to the model to obtain a classification label, then create a segmentation mask where the pixels within the ground truth polygon are assigned the predicted label, and compute class-weighted F1 between this mask and the ground truth mask. Finally, for SRE, we use the same evaluation scheme as building change detection.
@@ G Evaluation Setup
H Image Identifier Ablation
We run an ablation to measure the importance of including the image identifiers in the prompt on the temporal referring tasks (Table 9).
@@ H Image Identifier Ablation
Table 9: Impact of including image identifiers like ‘Image 1: ¡image¿ Image 2: ¡image¿‘ in the test-time prompt on TEOChat performance on the region-based temporal question answering and temporal referring expression tasks. Including image references leads to substantial benefits on both tasks.
I Single Image Qualitative Results
@@
We provide examples of responses output by TEOChat on the single image tasks (Figure 10).
@@
Table 10: Examples of responses produced by TEOChat on single image tasks.
@@
Automated methods for processing temporal EO data have the potential to lead to significant, positive societal impacts, from supporting disaster relief to monitoring deforestation. We show that TEOChat is capable of temporal reasoning, achieving strong performance on several real world tasks like building damage detection. We hope our work can inspire more work on developing vision and language assistants for important temporal EO tasks and lead to helpful new technologies that can benefit society.

### v2

@@ A Related Work
B Training Datasets
TEOChatlas consists of GeoChat_Instruct (Kuckreja et al., 2023), an instruction-following dataset with single earth observation images, as well as four temporal earth observation benchmarks (fMoW (Christie et al., 2018), xBD (Gupta et al., 2019), S2Looking (Shen et al., 2021), and QFabric (Verma et al., 2021)). We describe how we process each dataset in more detail below and a summary of the dataset statistics is provided in Table 7. We note that for all datasets, we (1) resize the shorter side to 224 then crop to 224x224 before inputting the images into the network and (2) randomly sample between all instruction-following tasks when creating examples, sometimes randomly sampling between different phrasings of the same task to introduce diversity in the wording. A spatial map of the locations of the examples in TEOChatlas is shown in Figure 5 and a comparison of TEOChatlas to other instruction-following EO datasets is shown in Table 6.
GeoChat_Instruct  The GeoChat_Instruct dataset (Kuckreja et al., 2023) is an instruction-following dataset composed of multiple single image EO datasets spanning tasks including object detection derived from SAMRS (which is itself a composition of DOTA (Xia et al., 2018), DIOR (Li et al., 2020), and FAIR1M (Sun et al., 2022)), visual question answering from LRBEN (Lobry et al., 2020) and Floodnet (Rahnemoonfar et al., 2021), and scene classification from NWPU-RESISC-45 (Cheng et al., 2017). The instruction-following tasks in the GeoChat_Instruct dataset include scene classification (classify the image into one of several classes provided in the prompt), visual question answering (answer a question about the image), referring expression (identify the location of an object given an expression describing its characteristics), region captioning (describe a location represented as a bounding box in the prompt), detailed description (describing the whole image in detail), grounded description (describing the whole image in detail with bounding boxes included in the description), and multi-round conversation (multiple turns of conversation). In total, the GeoChat_Instruct training dataset consists of 308,861 single instruction-following examples on 106,747 images.
@@ B Training Datasets
C Temporal Instruction-Following Tasks
@@
D Single Image Datasets
@@
E TEOChat Prompts
@@
A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. USER: This is a sequence of {high resolution, optical} satellite images {from <<<sensor>>>}: <<<video>>>.
where we randomly inject the information in curly brackets when metadata (resolution or sensor) is available and replace the “$<$video$>$” token with the “$<$image$>$” token $N$ times, with $N$ the number of images in the sequence. After this prompt we add the task instruction. Then, for tasks that require a choice of options, we provide the options in the prompt. For tasks that require bounding boxes to be output, we include “Please include bounding boxes of the form [x_min, y_min, x_max, y_max] in your response.” While this does not guarantee correct output parsing in every case, fewer than 0.1% errors occurred across all evaluations we conducted. Finally, we add “ASSISTANT:”.
F Additional Training Details
@@
G GeoChat Temporal Post-processing
@@
G.1 Temporal Scene Classification
@@ G.2 Change Detection
G.3 Spatial Referring Expression
@@
the spatS2Looking  The spatial referring expression tasks for S2Looking are to identify the constructed or destructed buildings. For both, we prompt GeoChat to identify all buildings in the first image and in the second image, construct a segmentation mask for each, then take a difference. We use the pixels in the first image but not the second as destructed and the inverse as constructed.
G.4 Change Question Answering
@@
G.5 Region-based Change Question Answering
@@
QFabric  These tasks are standard classification tasks, so we follow a similar procedure as for temporal scene classification and feed each image in the sequence (either 2 or 5 images) along with the bounding box in the prompt into GeoChat, then take a majority vote to get the final classification.
H Evaluation Setup
For previously existing tasks, we follow the metrics reported in the original works. For new tasks, we use accuracy. See further detail below.
Change Detection and Spatial Referring Expression  To evaluate TEOChat on building change detection (CD) and spatial referring expression (SRE), we follow the standard performance metrics of the building localization (Loc.) and damage classification (Dmg Cls.) subtasks on xBD and building change detection (Det.) subtask on S2Looking. Specifically, for building localization and building change detection, we convert the bounding box predictions and ground truth polygons to segmentation masks and compute per-pixel F1 between the predicted and ground truth segmentation masks. For building damage classification, we input the ground truth polygon to the model to obtain a classification label, then create a segmentation mask where the pixels within the ground truth polygon are assigned the predicted label, and compute class-weighted F1 between this mask and the ground truth mask. For SRE, we use the same evaluation scheme as building localization.
@@ H Evaluation Setup
I Additional Results
I.1 Image Reference Ablations
We run a ablation to investigate why image identifiers improve performance on the canonical bitemporal tasks (Table 9) and run another ablation to measure the importance of including the image identifiers in the prompt on the temporal referring tasks (Table 10).
Image Reference Strategy xBD Dmg Cls. (F1) S2Looking Det. (F1)
No References 44.7 25.5
Only ‘Image 1‘ 47.5 25.8
Only ‘Image 2‘ 49.2 34.5
Both ‘Image 1‘ and ‘Image 2‘ 50.0 34.5
Table 9: Impact of different image reference strategies on the canonical bitemporal datasets. Including ‘Image 2’ leads to much larger performance gains than including ‘Image 1‘, constituting most of the performance improvement from including both image references.
@@ I.1 Image Reference Ablations
Table 10: Impact of including image references in the test-time prompt on TEOChat’s performance on the region-based temporal question answering and temporal referring expression tasks. Including image references leads to substantial benefits on both tasks.
I.2 Fine-tuned Captioning Results
We evaluate TEOChat on three EO captioning datasets used as benchmarks in prior work (Hu et al., 2023; Zhan et al., 2024), namely UCM-Captions (Qu et al., 2016), Sydney-Captions Qu et al. (2016), and RSICD (Lu et al., 2017). Given these works include these datasets in the training data, we follow Hu et al. (2023) and fine-tune TEOChat on each of the three datasets separately. We use the same training procedure as used while training on TEOChatlas except we fine-tune until convergence (5, 20, and 5 epochs respectively). For comparison we also fine-tune GeoChat as results on these datasets are not reported in Kuckreja et al. (2023), and use their same training procedure when fine-tuning on each dataset for the same number of epochs as TEOChat. We evaluate using the same metrics reported in Hu et al. (2023) and Zhan et al. (2024). Fine-tuning results are reported in Tables 11, 12, 13 and 14 and zero-shot results are reported in Tables 15, 16, 17 and 18.
In the fine-tuning setting, we find TEOChat exceeds the performance of specialists and matches generalist single image EO models on the larger two of the three captioning datasets (UCM-captions and RSICD), and shows strong captioning abilities qualitatively on all three datasets (Figure 8). On Sydney-captions, TEOChat underperforms RSGPT, but qualitatively we find TEOChat’s captions are high quality. Importantly, TEOChat outperforms GeoChat across all metrics on all three captioning datasets except METEOR on UCM-captions, for which it slightly underperforms GeoChat (-0.4%). In the zero-shot setting, TEOChat performs similarly to or outperforms GeoChat on average across all metrics, with notable improvements on the largest, more difficult captioning dataset (RSICD). Interestingly, MiniGPT-v2 achieves impressive zero-shot performance on these datasets.
We believe the explanation for strong fine-tuning performance of RSGPT and MiniGPT-v2’s strong zero-shot and fine-tuning performance is twofold. First, we largely attribute this to the inclusion of a higher proportion of captioning datasets in their training data, whereas GeoChat and TEOChat have more scene classification and question answering tasks. Second, these three captioning datasets have several known issues which make their use as benchmarks suboptimal, including limited vocabulary and very short descriptions, and sometimes even misspelled words and grammatical errors. We caution researchers about using these benchmarks without careful analysis. We finally emphasize that in addition to these single image capabilities, TEOChat has temporal reasoning capabilities that none of these other models have.
Model BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROGUE_L CIDEr
Specialists
SAA(Lu et al., 2019) 79.6 74.0 69.1 64.8 38.6 69.4 294.5
Post-processing (Hoxha et al., 2023) 79.7 73.0 67.4 62.6 40.8 74.1 309.6
Generalists
MiniGPT-4 (Zhu et al., 2023b) 30.9 27.6 22.2 18.1 33.4 41.4 0.0
Shikra (Chen et al., 2023b) 81.2 58.9 43.3 34.0 32.6 56.7 56.7
MiniGPT-v2 (Chen et al., 2023a) 81.1 60.3 45.1 36.2 32.4 56.6 60.7
RSGPT (Hu et al., 2023) 86.1 79.1 72.3 65.7 42.2 78.3 333.2
GeoChat (Kuckreja et al., 2023) 82.9 74.9 68.5 63.0 44.4 78.1 299.1
TEOChat 85.2 78.2 72.4 67.4 43.9 79.3 316.5
Table 11: Captioning performance on the UCM-captions dataset.
Model BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROGUE_L CIDEr
Specialists
SAA (Lu et al., 2019) 68.8 60.7 52.9 45.4 30.5 58.2 170.5
Post-processing (Hoxha et al., 2023) 78.4 69.9 63.2 57.2 39.5 71.1 255.5
Generalists
MiniGPT-4 (Zhu et al., 2023b) 29.5 25.9 20.3 16.4 32.0 42.7 0.1
Shikra (Chen et al., 2023b) 77.5 53.2 37.0 27.8 29.4 53.3 26.8
MiniGPT-v2 (Chen et al., 2023a) 77.4 55.8 40.6 32.3 29.9 52.1 33.8
RSGPT (Hu et al., 2023) 82.3 75.3 68.6 62.2 41.4 74.8 273.1
GeoChat (Kuckreja et al., 2023) 74.1 61.2 53.1 47.0 34.2 67.8 184.3
TEOChat 75.1 64.6 56.9 50.5 36.2 69.5 204.6
Table 12: Captioning performance on the Sydney-captions dataset.
Model BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROGUE_L CIDEr
Specialists
SAA (Lu et al., 2019) 59.4 45.1 35.3 28.1 26.1 49.6 132.4
Post-processing (Hoxha et al., 2023) 62.9 46.0 35.7 28.7 25.3 47.3 75.6
Generalists
MiniGPT-4 (Zhu et al., 2023b) 34.0 31.8 25.8 20.6 33.2 40.7 0.1
Shikra (Chen et al., 2023b) 82.6 62.5 45.2 34.6 30.3 53.6 19.9
MiniGPT-v2 (Chen et al., 2023a) 83.1 64.6 47.8 37.3 30.2 54.0 52.4
RSGPT (Hu et al., 2023) 70.3 54.2 44.0 36.8 30.1 53.3 102.9
GeoChat (Kuckreja et al., 2023) 69.2 51.4 40.1 32.6 28.0 54.4 95.2
TEOChat 70.3 52.4 41.2 33.7 28.9 55.7 97.2
Table 13: Captioning performance on the RSICD captions dataset.
Model BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROGUE_L CIDEr
Specialists
SAA (Lu et al., 2019) 69.3 59.9 52.4 46.1 31.7 59.1 199.1
Post-processing (Hoxha et al., 2023) 73.7 63.0 55.4 49.5 35.2 64.2 213.6
Generalists
MiniGPT-4 (Zhu et al., 2023b) 31.5 28.4 22.8 18.4 32.9 41.6 0.1
Shikra (Chen et al., 2023b) 80.4 58.2 41.8 32.1 30.8 54.5 34.5
MiniGPT-v2 (Chen et al., 2023a) 80.5 60.2 44.5 35.3 30.8 54.2 49.0
RSGPT (Hu et al., 2023) 79.6 69.5 61.6 54.9 37.9 68.8 236.4
GeoChat (Kuckreja et al., 2023) 75.4 62.5 53.9 47.5 35.5 66.8 192.9
TEOChat 76.9 65.1 56.8 50.5 36.3 68.2 206.1
Table 14: Average performance across the three captioning datasets.
Model BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROGUE_L CIDEr
MiniGPT-v2 (Chen et al., 2023a) 17.6 9.1 4.2 1.9 14.8 23.5 16.9
GeoChat (Kuckreja et al., 2023) 22.8 10.3 4.3 1.7 10.9 21.5 13.9
TEOChat 8.8 4.3 1.9 0.8 10.3 20.8 15.7
Table 15: Zero-shot captioning performance on the UCM-captions dataset.
Model BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROGUE_L CIDEr
MiniGPT-v2 (Chen et al., 2023a) 18.6 8.1 2.8 1.0 13.5 24.4 10.0
GeoChat (Kuckreja et al., 2023) 19.8 6.9 0.0 0.0 8.5 19.1 4.3
TEOChat 28.4 8.4 2.1 0.0 8.8 21.8 4.1
Table 16: Zero-shot captioning performance on the Sydney-captions dataset.
Model BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROGUE_L CIDEr
MiniGPT-v2 (Chen et al., 2023a) 19.9 9.2 3.9 1.6 13.7 21.4 10.7
GeoChat (Kuckreja et al., 2023) 22.3 8.8 3.2 1.2 10.5 18.8 7.1
TEOChat 29.2 12.9 5.5 2.5 10.4 19.5 10.1
Table 17: Zero-shot captioning performance on the RSICD dataset.
Model BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROGUE_L CIDEr
MiniGPT-v2 (Chen et al., 2023a) 18.7 8.8 3.6 1.5 14.0 23.1 12.5
GeoChat (Kuckreja et al., 2023) 21.6 8.7 2.5 1.0 10.0 19.8 8.4
TEOChat 22.1 8.5 3.2 1.1 9.8 20.7 10.0
Table 18: Average zero-shot captioning performance across the three datasets.
I.3 Single Image Grounding Results
We evaluate TEOChat’s performance on single image grounding (Table 19). TEOChat outperforms MiniGPTv2 overall (+5.3%) and outperforms GeoChat in detecting large objects (+2.2%), but underperforms GeoChat across the other grounding tasks (-5.5% overall) which we suspect is due to the lower spatial resolution (224 vs. 504) required to handle temporal inputs. Qualitatively, TEOChat tends to make similar correct predictions and errors as GeoChat, but is a little less precise. To investigate the effect of lower spatial resolution on GeoChat’s performance, we evaluate GeoChat on 224x224 images (after performing bicubic interpolation of the positional embeddings to account for the change in spatial resolution). This results in huge drops (20-40% absolute drop) in grounding performance across all tasks, potentially suggesting a high sensitivity to input positional information on the grounding tasks.
Model Small Medium Large Single-object grounding Multi-object grounding [refer] [grounding] Overall
MiniGPT-v2 (Chen et al., 2023a) 14.3 37.3 54.7 38.0 13.1 33.1 18.3 32.0
GeoChat (Kuckreja et al., 2023) 26.4 50.4 63.1 50.6 22.8 45.1 27.5 43.8
GeoChat 224 x 224 1.1 4.0 16.3 6.3 2.3 5.4 4.6 5.3
TEOChat 16.3 43.0 65.5 43.1 19.1 38.2 26.2 37.3
Table 19: Single EO image grounding performance (Acc@0.5). We evaluate all tasks using axis-aligned bounding boxes. All metrics are defined as in Kuckreja et al. (2023).
I.4 Temporal Sequence Sampling Ablation
We run an ablation to measure the impact of the random samples of 8 images taken when the image sequences are longer than 8 images on fMoW RGB and fMoW Sentinel (Table 20).
Run fMoW RGB fMoW Sentinel
1 75.1 45.5
2 75.2 45.3
3 75.1 45.6
4 75.1 45.4
5 75.2 45.4
Table 20: Impact of random samples of 8 images from long image sequences on fMoW RGB and fMoW Sentinel validation set performace. There is minimal variance (0.1-0.2%) from different random samples.
I.5 Additional Temporal Qualitative Results
@@
Examples of TEOChat’s responses on canonical temporal tasks are shown in Figure 6.
I.6 Single Image Qualitative Results
@@
Examples of responses output by TEOChat on various single EO image tasks are shown in Figure 7, and examples of TEOChat responses on single EO image captioning tasks are provided in Figure 8.
Figure 7: Examples of TEOChat’s responses on single image tasks.
Figure 8: Examples of TEOChat’s responses on the three single image EO captioning datasets after fine-tuning following Hu et al. (2023).
@@
We show that TEOChat is capable of temporal reasoning, achieving strong performance on several real world tasks like building damage detection. Automated, generalist methods for processing temporal EO data like TEOChat have the potential to lead to significant, positive societal impacts. For example, we believe that such methods could be deployed at scale on sequences of Earth observation data to automatically derive information, ranging from disaster response to urban development monitoring. This analysis is typically conducted by experts acquiring data on the ground or by manually inspecting UAV/satellite imagery, which becomes impractical to do in large areas which may require several thousands of high resolution images to capture. Such technology could, for example, aid government agencies and humanitarian organizations to acquire information much more rapidly, which is key for disaster response (Khan et al., 2023). We hope our work can inspire more work on developing vision and language assistants for important temporal EO tasks and lead to helpful new technologies that can benefit society.